OpenStudy (mtalhahassan2):

Nina deposits $3500 into a savings account. The rate of simple interest is 5.5%/a. a) By how much does the amount in her account increase each year? b) Determine the amount in her account at the end of each of the first 5 years. c) State the total amount as the general term of a sequence. d) Graph this sequence

OpenStudy (danjs):

Do you remember the formula for simple interest SI from class

OpenStudy (mtalhahassan2):

yes I do

OpenStudy (danjs):

what is that

OpenStudy (mtalhahassan2):

I=Prt

OpenStudy (danjs):

ok, what is the principle P

OpenStudy (mtalhahassan2):

3500

OpenStudy (danjs):

the rate is 5.5% or 5.5/100 or 0.055

OpenStudy (mtalhahassan2):

yep

OpenStudy (mtalhahassan2):

I geted the a

OpenStudy (mtalhahassan2):

192.5

OpenStudy (danjs):

Interest = 3500 * 0.055 * period What is the period for the interest

OpenStudy (danjs):

Interest = 3500 * 0.055*t t = # years right?

OpenStudy (mtalhahassan2):

for the part b

OpenStudy (danjs):

so at the end of the first year, what is the total value of the account

OpenStudy (danjs):

Principle + interest

OpenStudy (mtalhahassan2):

oh ok

OpenStudy (danjs):

That will be your principle for the second year,

OpenStudy (mtalhahassan2):

3692.5

OpenStudy (mtalhahassan2):

for the 1 year

OpenStudy (danjs):

actually , no this will be just simple interest, a straight line, the increase in the account value will not be part of the next interest calculation right

OpenStudy (danjs):

You always use the same principle 3500 for any term length... like for 5 years it is still Interest = 3500 * 0.055 * 5

OpenStudy (mtalhahassan2):

o it be like 962.5

OpenStudy (danjs):

yeah, add that to the principle, that is the amount in the account at the end of 5 years

OpenStudy (mtalhahassan2):

but they said we have to find for the whole year like from 1 years to 5 years

OpenStudy (mtalhahassan2):

i find the first year year and it be 369.50

OpenStudy (danjs):

Total amount in her account after t years will be Total Account = 3500 + I(t) where I(t) = 3500*0.055*t

OpenStudy (mtalhahassan2):

know how can we find the amount for the 2 year

OpenStudy (danjs):

use t = 2, 3, 4

OpenStudy (danjs):

Year 2 Total = 3500 + I(2) I(2) = 3500*0.055*2 = 385 Total = 3500 + 385 = 3885

OpenStudy (danjs):

Same thing for t = 3 and t=4 and t = 5

OpenStudy (danjs):

To graph it, you want the Y axis to be the total money in the account, and the X axis to be time t in years. Total in account = Principle + interest y = 3500 + 3500*0.05*x

OpenStudy (mtalhahassan2):

why u write 3500 two times

OpenStudy (mtalhahassan2):

and what about part c

OpenStudy (danjs):

Because you want the total amount of money in the account, which is the principle 3500, plus the interest, which is 3500*0.55*t

OpenStudy (mtalhahassan2):

what about part c

OpenStudy (danjs):

Part c,i think just wants you to write the general function for the value of the account after t years. Which is, Total = Principle + interest Total = 3500 + 3500*0.55*t or factored Total = 3500*( 1 + 0.055 * t )

OpenStudy (danjs):

here is the graph...

OpenStudy (danjs):

notice at the start, x=0, the y axis is the initial value 3500 dollars,

OpenStudy (mtalhahassan2):

for the part c the answer given in my book is like tn= 3500+192.50

OpenStudy (danjs):

right, that is the same as tn = Principle + interest tn = 3500 + 3500*0.055*n for n=1,2,3....

OpenStudy (danjs):

tn - total value of account at n year

OpenStudy (mtalhahassan2):

tn = 3500*0.055

OpenStudy (mtalhahassan2):

192.5

OpenStudy (danjs):

Yes the overall general function is Total = Principle + interest = 3500 + 192.5*t

OpenStudy (mtalhahassan2):

ok but where u get 3500 again

OpenStudy (danjs):

that is the principle amount she deposited into the account 3500 dollars The interest at any time t in years is, 3500*0.055*t = 192.5*t So the total amount in the account after a time t years is the sum of the initial principle 3500 and the interest 192.5*t Total = 3500 * 192.5*t

OpenStudy (danjs):

If you want to graph that, let y = total amount x = time t in years y = 3500 + 192.5x

OpenStudy (mtalhahassan2):

ok thanks a lot

OpenStudy (danjs):

welcome, anytime